THE NUMBER OF SOLUTIONS OF THE EQUATION


Z^2 + Z' = 0Ā  IS -------


A) 4


B)3


C)2


D)1

Grade:12

1 Answers

Vikas TU
10 years ago

Put z = x + iy

and z'' = x - iy

 

(x + iy)^2 + x - iy = 0

x^2 - y^2 + 2xyi + x - iy = 0

(x^2 - y^2 + x) + i(2xy - y) = 0

 

1) 2xy - y = 0

    y(2x - 1) = 0

 y = 0 , x = 1/2

 

2) x^2 - y^2 + x= 0

    at x = 1/2 y= +root3/2, -root3/2

    at y = 0 x = 0, -1

 

Therefore answer shoulbe (a) 4 solutions.

 

 

PLZ  APPROVE 

Dear frnd!

 

 

2) x^2 - y^2 + x= 0